y is directly proportional to x^2 and y=a for a particular value of x. Find an expression for y in terms of a, when this value of x is doubled.

Please help me with this!! Thank you.

Answer by akch2002(12)

y is directly proportional to x^2 and y=a for a particular value of x. Find an expression for y in terms of a, when this value of x is doubled.
we can write
y =kx^2 where k is proportionality constant.
further
a = kx^2 ---- (1)
We have to find when x is doubled.
Let it be y1.
Hence, y1= k(2x)^2 = k*4x^2 ---- (2)
Putting value of k from 1, we get
y1 = (a/x^2)*4x^2
= 4a will be the value of when at double value.